OpenStudy (anonymous):

The cost of manufacturing x cases of fillfluff is C dollars (in thousands), where C(x) = 4 x + 4 sqrt(x) + 5 Monthly production at t weeks is expected to be x = 6400 + 90 t Write an equation to express rate of change of cost with respect to time: (dC)/(dt) = i found equation to express manufacturing cost as a function of time: C(t) =25600+360t+4sqrt(6400+90t)+5 and the equation to express marginal cost: (dC)/(dx) =4+2/sqrt(x)

OpenStudy (anonymous):

Chain rule. \[dC/dt=dC/dx*dx/dt\]

OpenStudy (anonymous):

Another thing to consider is changing to exponential notation for the square root. Then you can use the exponential rule for taking the derivative (easier I think).

OpenStudy (anonymous):

Having a tough time tonite. Here it is....\[dC/dt=dC/dx∗dx/dt=(4+2(6400+90t)^{-1/2})∗90\]

OpenStudy (anonymous):

Note that \[x=x(t)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so did you see what i found?

OpenStudy (anonymous):

so does the rate of change mean take the derivitive of the marginal cost?

OpenStudy (anonymous):

can that be reduced any more ?